find the exact distance between the points
(12,20) and (2,8)

Answer by Alan3354(30993)

find the exact distance between the points
(12,20) and (2,8)
----------
s^2 = diffy^2 + diffx^2 = 12^2 + 10^2
s^2 = 244
s = 2sqrt(61)